# CosyVoice 2: Scalable Streaming Speech Synthesis with Large Language Models

### 13 Dec 2024 v1

@@ 3.2 Training Data for CosyVoice 2
CosyVoice 2 shares the same training data as its previous version [34]. We first collect the speech-only data with internal speech processing tools, including speech detection, signal-to-noise ratio (SNR) estimation, speaker diarization, and separation. Subsequently, the Paraformer [50] and SenseVoice [43] are employed to generate pseudo text labels for Chinese and other languages, respectively. We also employ an internal force-alignment model to filter out low-quality data and enhances the accuracy of punctuation. Data details are provided in Table 3.
@@ 3.3 Evaluation Settings
We evaluate our CosyVoice 2 on two test sets. The first one is constructed from the test-clean set of Librispeech corpus [51], denoting as test-clean. This test set is used to evaluate CosyVoice 2 on a limited English domain. The Whisper-large V3 is used as the ASR model to evaluate the content consistency. As for the speaker similarity (SS), we employ the ERes2Net model [52] to extract speaker embeddings of prompt and generated utterances, and their raw cosine similarity is treated as the speaker similarity. NMOS score [53] is used to evaluate the objective quality.
The second evaluation is conducted under the SEED settings [33], which is widely used to evaluate recent TTS models, covering various text domains and reference speeches. In this evaluation, about 2,000 Chinese and 1,000 English samples are selected from CommonVoice datasets, denoting as test-zh and test-en, respectively. In addition, about 400 hard test cases are also included to evaluate the robustness of TTS models on text repetition, tongue twister and other challenging synthesis cases, denoting as test-hard in this report. The Paraformer is employed to recognize the synthesis results of test-zh and test-hard, while the Whisper-large V3 is adopted for test-en to evaluate the content consistency.
@@ 4.2 Comparison Results with Baselines
CER (%) SS WER (%) SS WER (%) SS
Human 1.26 0.760 2.14 0.730 - -
@@
FireRedTTS [35] 1.51 0.630 3.82 0.460 17.45 0.639
MaskGCT [18] 2.27 0.774 2.62 0.774 10.27 0.748
E2 TTS (32 NFE) [31] 1.97 0.73 2.19 0.710 - -
F5-TTS (32 NFE) [32] 1.56 0.76 1.83 0.670 8.67 0.762
CosyVoice [34] 3.63 0.775 4.29 0.699 11.75 0.755
CosyVoice 2 1.45 0.806 2.57 0.736 6.83 0.776
CosyVoice 2-S 1.45 0.812 2.38 0.743 8.08 0.785
Table 6: Results of CosyVoice 2 and recent TTS models under the SEED evaluation settings. $\dagger$ denotes close-sourced models.
We also evaluated CosyVoice 2 on the commonly-used test sets: SEED test-zh, test-en and test-hard, which include diverse input texts and reference speeches from various domains. The experimental results for CosyVoice 2 and the baseline models are presented in Table 6. On the test-zh set, CosyVoice 2 surpasses all open-sourced models in terms of CER and SS, falling short of the commercial model SEED-TTS by only a small margin. On the test-en set, CosyVoice 2 ranks fourth and third in terms of WER and SS, respectively. This may result from the imbalance in the volume of training data between Chinese and English. We plan to explore data scaling in future work to enhance content consistency in English. On the test-hard set, the offline CosyVoice 2 model achieves state-of-the-art performance across all compared baseline, demonstrating its robustness in challenging synthesis scenarios. Compared with human-generated speeches, CosyVoice 2 shows comparable content consistency and superior speaker similarity. Considering the recognition errors can also stem from the ASR model, it is reasonable to conclude that CosyVoice 2 achieves a human-parity synthesis capability. We also evaluated the streaming mode, denoted as “CosyVoice 2-S” in Table 5 and 6. For both evaluation settings, the streaming mode’s performance is nearly lossless in typical test cases. Only in challenging cases is there a slight degradation in content consistency, highlighting the strength of our unified streaming/non-streaming framework.
@@ 4.5 Results on Instructed Generation
CosyVoice-Instruct [34] 1.72 0.797 3.94 3.14
CosyVoice 2 1.52 0.804 3.94 4.11
CosyVoice 2 w/o Instruction 0.97 0.817 4.02 2.54
@@ 4.7 LM Fine-tuning with Reinforcement Learning
Although the SFT can improve the performance on most speakers, the results of Spk E are still worse than the base model especially on English. Because it has a more complex voice and faster speech speed. Additionally, only Chinese recordings are available for Spk E. So we choose the Spk E to evaluate the effectiveness of reinforcement learning. During the SFT of text-speech LM, we apply the DPO to change the preference biasing of the LM by the ASR and SS rewards (DPO-ASR-SS). We also use the differentiable ASR rewards to optimize the LM parameters. After SFT, we evaluate the model with content consistency (WER), speaker similarity (SS) and speech quality (NMOS) on the test set of Spk E and further evaluated the WER on the SeedTTS test sets to explore whether the model can maintain robustness to out-of-domain or cross-lingual input text. Results are shown in Table 11.

### 18 Dec 2024 v2

@@ 3.2 Training Data for CosyVoice 2
CosyVoice 2 shares the same training data as its previous version [34]. We first collect the speech-only data with internal speech processing tools. Subsequently, the Paraformer [50] and SenseVoice [43] are employed to generate pseudo text labels for Chinese and other languages, respectively. We also employ an internal force-alignment model to filter out low-quality data and enhances the accuracy of punctuation. Data details are provided in Table 3.
@@ 3.3 Evaluation Settings
We evaluate our CosyVoice 2 on two test sets. The first one is constructed from the test-clean set of Librispeech corpus [51], denoted as test-clean. This test set is used to evaluate CosyVoice 2 on a limited English domain. The Whisper-large V3 is used as the ASR model to evaluate the content consistency. As for the speaker similarity (SS), we employ the ERes2Net model [52] to extract speaker embeddings of prompt and generated utterances, and their raw cosine similarity is treated as the speaker similarity. NMOS score [53] is used to evaluate the objective quality.
The second evaluation is conducted on the SEED test sets [33], which is widely used to evaluate recent TTS models, covering various text domains and reference speeches. In this evaluation, about 2,000 Chinese and 1,000 English samples are selected from CommonVoice datasets, denoted as test-zh and test-en, respectively. In addition, about 400 hard test cases are also included to evaluate the robustness of TTS models on text repetition, tongue twister and other challenging synthesis cases, denoted as test-hard in this report. The Paraformer is employed to recognize the synthesis results of test-zh and test-hard, while the Whisper-large V3 is adopted for test-en to evaluate the content consistency. We adopt two speaker verification (SV) models to evaluate speaker similarity: WavLM-finetuned SV model and ERes2Net.
@@ 4.2 Comparison Results with Baselines
CER (%) $\downarrow$ SS $\uparrow$ WER (%) $\downarrow$ SS $\uparrow$ WER (%) $\downarrow$ SS $\uparrow$
Human 1.26 0.755 (0.775) 2.14 0.734 (0.742) - -
@@
FireRedTTS [35] 1.51 0.635 (0.653) 3.82 0.460 (0.526) 17.45 0.621 (0.639)
MaskGCT [18] 2.27 0.774 (0.752) 2.62 0.714 (0.730) 10.27 0.748 (0.720)
E2 TTS (32 NFE) [31] 1.97 0.730 2.19 0.710 - -
F5-TTS (32 NFE) [32] 1.56 0.741 (0.794) 1.83 0.647 (0.742) 8.67 0.713 (0.762)
CosyVoice [34] 3.63 0.723 (0.775) 4.29 0.609 (0.699) 11.75 0.709 (0.755)
CosyVoice 2 1.45 0.748 (0.806) 2.57 0.652 (0.736) 6.83 0.724 (0.776)
CosyVoice 2-S 1.45 0.753 (0.812) 2.38 0.654 (0.743) 8.08 0.732 (0.785)
Table 6: Results of CosyVoice 2 and recent TTS models on the SEED test sets. $\dagger$ denotes close-sourced models. For speaker similarity, the result in a bracket are measured by ERes2Net, while the results outside brackets are measured by WavLM-based models.
We also evaluated CosyVoice 2 on the commonly-used test sets: SEED test-zh, test-en and test-hard, which include diverse input texts and reference speeches from various domains. The experimental results for CosyVoice 2 and the baseline models are presented in Table 6. On the test-zh set, CosyVoice 2 surpasses all open-sourced models in terms of CER and SS, falling short of the commercial model SEED-TTS by only a small margin. On the test-en set, CosyVoice 2 ranks fourth and third in terms of WER and SS, respectively. This may result from the imbalance in the volume of training data between Chinese and English. We plan to explore data scaling in future work to enhance content consistency in English. On the test-hard set, the offline CosyVoice 2 model achieves state-of-the-art performance across all compared baseline, demonstrating its robustness in challenging synthesis scenarios. Compared with human-generated speeches, CosyVoice 2 shows comparable content consistency and superior speaker similarity. Considering the recognition errors can also stem from the ASR model, it is reasonable to conclude that CosyVoice 2 achieves a human-parity synthesis capability. We also evaluated the streaming mode, denoted as “CosyVoice 2-S” in Table 5 and 6. For both evaluation settings, the streaming mode’s performance is nearly lossless in typical test cases. Only in challenging cases is there a slight degradation in content consistency, highlighting the strength of our unified streaming/non-streaming framework. We found that the results of speaker similarity are not consistent on different SV models. This may indicate a new research topic on how to evaluate speaker similarity for TTS models automatically. Since different TTS models may use different SV models to extract speaker information, evaluating speaker similarity with the same SV model allows a more accurate evaluation on the utilization of speaker information. Therefore, we employ ERes2Net  [^2] https://github.com/modelscope/3D-Speaker for evaluating speaker similarity in subsequent experiments.
@@ 4.5 Results on Instructed Generation
CosyVoice-Instruct [34] 1.72 0.797 3.94 3.09
CosyVoice 2 1.52 0.804 3.94 4.06
CosyVoice 2 w/o Instruction 0.97 0.817 4.02 2.28
@@ 4.7 LM Fine-tuning with Reinforcement Learning
Although the SFT can improve the performance on most speakers, the results of Spk E are still worse than the base model especially on English. Because Spk E has a more complex voice and faster speech speed. Additionally, only Chinese recordings are available for Spk E. So we apply the reinforcement learning on Spk E for further improvements. For DPO, we synthesis 10 thousand samples-pairs by the SFT models to change the preference biasing of the LM by the ASR and SS rewards. We also use the differentiable ASR rewards to optimize the LM parameters. After RL, we evaluate the model with content consistency (WER), speaker similarity (SS) and speech quality (NMOS) on the test set of Spk E and further evaluated the WER on the SeedTTS test sets to explore whether the model can maintain robustness to out-of-domain or cross-lingual input text. Results are shown in Table 11.